### v1

@@ 3 Modeling Restricted Decoding Problems
Overall, we obtain the system of polynomial constraints $F(\mathbf{x})=(f_{1},\ldots,f_{n})\in\mathbb{F}[x_{1},\ldots,x_{k}]^{n}$ with
@@
$f_{i}(\mathbf{x})=f_{\mathcal{E}}\left(x_{i}\right)\quad\text{for }i\in[k],\quad\text{and}\quad f_{k+i}(\mathbf{x})=f_{\mathcal{E}}\left(s_{i}-\langle\mathbf{a}_{i},\mathbf{x}\rangle\right)\quad\text{for }i\in[r],$ (1)
@@
where $\mathbf{a}_{i}\in\mathbb{F}^{k}$ denotes the $i$-th row of $\mathbf{A}$. The correctness and degree of this modeling are stated in the following proposition.
Proposition 1
@@
The polynomial relation $F$ given in Equation 1 provides a degree-$z$ modeling of $\mathcal{E}$-SDP: If $\mathbf{w}\in\mathbb{F}^{k}$ satisfies $F\left(\mathbf{w}\right)=\mathbf{0}$, then the error vector $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ solves the $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ with $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$.
Proof
@@
3.0.4 Performance in TCitH framework.
From the provided modeling of $\mathcal{E}$-SDP, a digital signature is derived using the TCitH framework [21], as summarized in Section 2.0.4. The chosen parameters and the obtained signature sizes are provided in Table 1. The parameters $\tau$, $N$, and $w$ of the proof system are selected following [14] to allow for comparable runtime performance of the VC scheme. To simplify implementation, the smallest extension degree $\mu=[\mathbb{K}:\mathbb{F}]$ satisfying the requirement $N\leq\left\lvert\mathbb{K}\right\rvert=\left\lvert\mathbb{F}\right\rvert^{\mu}$ is chosen. This, in turn, determines the batching parameter $\eta$.
Table 1: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the TCitH framework. The proof system parameters are selected analogously to those in MQOM [14] to ensure comparable runtime performance of the VC scheme.
Security
category
$\mathcal{E}$-SDP Proof System Performance
@@
NIST 1
$127$
77
$127$
$76$
$24$ $256$ 22 $10$ 44 fast $7650$
$15$ $2048$ 22 $10$ 66 short $5533$
33
22
$579$
$213$
$17$ $256$ 66 $14$ 99 fast $3533$
$12$ $2048$ 77 $12$ 88 short $3095$
NIST 3
$127$
77
$187$
$111$
$36$ $256$ 22 $14$ 66 fast 16 67516\,675
$23$ $2048$ 22 $14$ 44 short 12 35412\,354
33
22
$839$
$309$
$27$ $256$ 66 2121 33 fast $8284$
$18$ $2048$ 77 $18$ $12$ short $6860$
NIST 5
$127$
77
$251$
$150$
$48$ $256$ 22 1919 77 fast 29 83929\,839
$31$ $2048$ 22 1919 33 short 22 30522\,305
33
22
$1102$
$406$
$36$ $256$ 66 $27$ 44 fast 14 58414\,584
$25$ $2048$ 77 $24$ 66 short 12 60812\,608
@@
3.0.5 Performance in VOLEitH framework.
Table 2 provides an overview of the performances that can be achieved if TCitH is replaced by the VOLEitH framework. Parameters of the proof system are chosen analogously to SDitH [3]. For Ternary-SDP, only a minor improvement in signature size is achievable. A similar effect is observed in [14]: for OWFs with small degree and small witness size, the structurally simpler TCitH framework remains competitive. On the other hand, for CROSS-SDP, which has a degree of $d=z=7$, VOLEitH enables a reduction in signature size of about $1\text{\,}\mathrm{kB}$ for NIST security category 1, and of up to $6\text{\,}\mathrm{kB}$ for NIST security category 5.
Table 2: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the VOLEitH framework. The proof system parameters are selected analogously to those in SDitH [3] to ensure comparable runtime performance of the VC scheme.
Security
category
$\mathcal{E}$-SDP Proof System Performance
[…]
NIST 1
$127$
77
$127$
$76$
$16$ $256$ $32$ $101$ 33 fast $6432$
$11$ $2048$ $22$ $107$ $10$ short $4372$
33
22
$579$
$213$
$16$ $256$ $96$ $101$ 11 fast $3515$
$11$ $2048$ $81$ $107$ 88 short $2974$
NIST 3
$127$
77
$187$
$111$
$24$ $256$ $48$ $153$ 33 fast 14 35914\,359
$16$ $4096$ $32$ $157$ 33 short $9361$
33
22
$839$
$309$
$24$ $256$ $144$ $153$ 11 fast $7816$
$16$ $4096$ $128$ $157$ 11 short $6463$
NIST 5
$127$
77
$251$
$150$
$32$ $256$ $64$ $206$ 33 fast 25 57325\,573
2121 $4096$ $42$ $216$ 77 short 16 58916\,589
33
22
$1102$
$406$
$32$ $256$ $192$ $206$ 11 fast 13 85113\,851
2121 $4096$ $168$ $216$ 55 short 11 52111\,521
3.0.6 Comparison with NIST submissions.
A comparison of the achieved signature sizes with submissions to NIST’s call for additional digital signature schemes [24] is provided in Table 3. As can be seen, the TCitH and VOLEitH frameworks can be combined with a wide range of OWFs. The proposed modeling of $\mathcal{E}$-SDP is competitive compared to the state-of-the-art: for the Ternary-SDP, we obtain a signature size which is very close to those of the smallest proposals that utilize the TCitH or the VOLEitH framework with an alternative hardness assumption. The sizes obtained from CROSS-SDP are slightly larger than those of SDitH and PERK, which also use modelings of degree greater than two, but smaller than $z=7$. The code-based proposals CROSS and LESS target other trade-offs: CROSS uses a simple CVE-like protocol [18, 6], and LESS achieves a small signature size by using multiple public keys (resulting in a public-key size that is multiple times larger than those of the other schemes shown).
Table 3: Comparison of the obtained signature sizes with NIST submissions. The shown sizes correspond to the short optimization tradeoff; runtime may vary.
Scheme Hardness Assumption Design Principle $\left\lvert\sigma\right\rvert$ in $\mathrm{kB}$]
@@
4 Conclusion

### v2

@@ II TCitH and VOLEitH frameworks
NIST 1 $127$ 77 $127$ $76$ $24$ $256$ 22 $10$ 44 fast $7650$
$15$ $2048$ 22 $10$ 66 short $5533$
33 22 $579$ $213$ $17$ $256$ 66 $14$ 99 fast $3533$
$12$ $2048$ 77 $12$ 88 short $3095$
NIST 3 $127$ 77 $187$ $111$ $36$ $256$ 22 $14$ 66 fast 16 67516\,675
$23$ $2048$ 22 $14$ 44 short 12 35412\,354
33 22 $839$ $309$ $27$ $256$ 66 2121 33 fast $8284$
$18$ $2048$ 77 $18$ $12$ short $6860$
NIST 5 $127$ 77 $251$ $150$ $48$ $256$ 22 1919 77 fast 29 83929\,839
$31$ $2048$ 22 1919 33 short 22 30522\,305
33 22 $1102$ $406$ $36$ $256$ 66 $27$ 44 fast 14 58414\,584
$25$ $2048$ 77 $24$ 66 short 12 60812\,608
III Modeling Restricted Decoding Problems
We now present a polynomial modeling for $\mathcal{E}$-SDP that can be utilized in the TCitH and VOLEitH frameworks. First, we recap some basic notions from coding theory.
Coding theory. A linear code over $\mathbb{F}_{p}$ of length $n$ is a $k$-dimensional subspace of $\mathbb{F}_{p}^{n}$. The code can be represented by a parity-check matrix $\mathbf{H}\in\mathbb{F}_{p}^{(n-k)\times n}$ such that $\mathbf{c}\mathbf{H}^{\top}=\mathbf{0}$ for every codeword $\mathbf{c}$. We denote by r=nkr=n-k the redundancy of the code. The syndrome of a vector $\mathbf{v}$ (through $\mathbf{H}$) is $\mathbf{s}=\mathbf{v}\mathbf{H}^{\top}\in\mathbb{F}_{p}^{r}$. The Hamming weight (from now on, only weight) of a vector $\mathbf{v}$ is the number of non-zero entries in $\mathbf{v}$.
Restricted decoding assumption. The (general) restricted Syndrome Decoding Problem ($\mathcal{E}$-SDP) [19] replaces the Hamming-weight constraint of the usual SDP with the requirement that each entry must lie in the fixed subset $\mathcal{E}$ of the underlying finite field $\mathbb{F}_{p}$. It is defined as follows:
Definition 2 ($\mathcal{E}$-SDP).
Let the restriction $\mathcal{E}\subset\mathbb{F}_{p}$, the parity-check matrix $\mathbf{H}\in\mathbb{F}_{p}^{r\times n}$, and the syndrome $\mathbf{s}\in\mathbb{F}_{p}^{r}$ be given. Solving $\mathcal{E}$-SDP requires to find $\mathbf{e}\in\mathcal{E}^{n}$ such that $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$.
We refer to the size of the restriction as $z=\left\lvert\mathcal{E}\right\rvert$. Different choices of $\mathcal{E}$ are possible, and all of them lead to an NP-complete problem [20]. In this work, we mainly focus on two specific choices that have undergone careful analysis.
CROSS restriction
The signature scheme CROSS [12] uses the restriction $\mathcal{E}=\{2^{i}\mid i\in[7]\}$ within $\mathbb{F}_{p}$ of size $p=127$. We are going to refer to this setting as CROSS-SDP, the concrete hardness of which is analyzed in detail in [21] and confirmed in [22]. For instance, the parameters $n=127$ and $k=76$ provide a suitable choice for NIST security category 1.
Ternary decoding with full weight
The signature scheme WAVE [14, 13] relies on the ternary syndrome decoding with high weight. For state-of-the-art solvers [23], parameterizing the problem such that the error is full-weight and unique, maximizes the concrete hardness (a setting in which WAVE cannot sign successfully). This corresponds to $\mathcal{E}$-SDP with $\mathcal{E}=\{1,2\}$ and $p=3$, to which we are going to refer as Ternary-SDP. When polynomial factors are conservatively ignored, the solver presented in [23] requires $2^{0.247\cdot n}$ operations for $k\approx 0.369\cdot n$, which implies uniqueness of the solution. In particular, the parameters $n=579$ and $k=213$ are shown to achieve NIST security category 1.
General $\mathcal{E}$-SDP
In general, any choice of restriction requires a careful analysis of the problem’s concrete computational hardness. This makes suggesting parameters difficult if one deviates from the settings described above. Nevertheless, we are interested in assessing whether other parameterizations can yield an improved performance in combination with the proposed modeling. An upper bound on the cost of solving a general $\mathcal{E}$-SDP instance is provided by Stern’s algorithm, as adapted in [20]. The following proposition provides a tight closed-form approximation for the cost of this solver.
Proposition 1.
Let $Z=\log_{2}\left(z\right)$, $P=\log_{2}(p)$, and k/n=1Z/Pk/n=1-Z/P. Then, conservatively ignoring polynomial factors and memory cost, Stern’s solver has time complexity $C$ with
$\log_{2}\left(C\right)\approx\frac{P-Z}{2\cdot P-Z}\cdot Z\cdot n.$ (1)
Proof.
For k/n=1Z/Pk/n=1-Z/P, the solution $\mathbf{e}$ is unique with good probability [21]. Following [20], the cost of Stern’s solver is
$\min_{0\leq\ell\leq r}\left\{z^{\frac{k+\ell}{2}}+\frac{z^{k+\ell}}{p^{\ell}}\right\},$
when polynomial factors and memory cost are neglected. The optimal $\ell$ balances $z^{\frac{k+\ell}{2}}=\frac{z^{k+\ell}}{p^{\ell}}$; this is achieved for $k/\ell=2P/Z-1$. Equation˜1 follows from $C\approx z^{\frac{k+\ell}{2}}=p^{\ell}$. ∎
Progress in cryptanalysis of $\mathcal{E}$-SDP would require increased $n$. As in the case of [24], since the signature size is typically dominated by $\sigma_{\mathrm{sym}}$ for the modeling we propose, this would have only a minor impact on performance in the TCitH and VOLEitH frameworks.
Modeling $\mathcal{E}$-SDP. The linearity of the syndrome equations allows a standard optimization that reduces the witness size from $n$ to $k$. We assume that the parity-check matrix is in systematic form  [^5] This assumption does not impact security. $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$, where $\mathbf{A}\in\mathbb{F}_{p}^{r\times k}$ and $\mathbf{1}_{r}\in\mathbb{F}_{p}^{r\times r}$ denotes the identity matrix. Then, $\mathbf{e}\in\mathbb{F}_{p}^{n}$ satisfies $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$ if and only if it is of the form $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ with $\mathbf{w}\in\mathbb{F}_{p}^{k}$. That is, the witness is defined as a length-$k$ partial error vector that is expanded to a length-$n$ error vector so that the syndrome equations are satisfied.
[…]
@@ III Modeling Restricted Decoding Problems
Overall, we obtain the system of polynomial constraints $F(\mathbf{x})=(f_{1},\ldots,f_{n})\in\mathbb{F}_{p}[x_{1},\ldots,x_{k}]^{n}$ with
$f_{i}(\mathbf{x})$ $=f_{\mathcal{E}}\left(x_{i}\right)\quad\text{for }i\in[k],$
@@
$\text{and}\quad f_{k+i}(\mathbf{x})$ $=f_{\mathcal{E}}\left(s_{i}-\langle\mathbf{a}_{i},\mathbf{x}\rangle\right)\quad\text{for }i\in[r],$ (2)
@@
where $\mathbf{a}_{i}\in\mathbb{F}_{p}^{k}$ denotes the $i$-th row of $\mathbf{A}$. The correctness and degree of this modeling are stated in the following proposition.
Proposition 2.
@@
The polynomial relation $F$ given in Equation˜2 provides a degree-$z$ modeling of $\mathcal{E}$-SDP: $\mathbf{w}\in\mathbb{F}_{p}^{k}$ satisfies $F\left(\mathbf{w}\right)=\mathbf{0}$, if and only if the error vector $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ solves the $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ with $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$.
Proof.
@@
$\mathcal{E}$-SDP-enabled tradeoffs. For the proposed modeling, the signature sizes in both TCitH and VOLEitH scale as $\left\lvert\sigma_{F}\right\rvert\propto z$ and $\left\lvert\sigma_{\mathbf{w}}\right\rvert\propto k\cdot\log_{2}(p)$, while $\sigma_{\mathrm{sym}}$, composed of symmetric primitives, is largely independent of the modeling. Figure˜2 shows the tradeoff between $k\cdot\log_{2}(p)$ and $z$ which can be achieved by different parameterizations of $\mathcal{E}$-SDP. The solver cost is approximated by Proposition˜1. This approximation is tight for both Ternary-SDP and CROSS-SDP, parametrized according to the careful analysis in [23] and [21], respectively. We observe that only a minor improvement in witness size may be possible by increasing the degree of the relation (i.e., the size of $\mathcal{E}$). For fixed $z$, smaller field sizes $p$ are preferable. In the following, we focus on CROSS-SDP and Ternary-SDP; by Figure˜2, the latter is expected to yield the smallest $\mathcal{E}$-SDP-based signatures.
@@
Figure 2: Tradeoff between degree and witness size. Black lines correspond to $\mathcal{E}$-SDP parameters achieving NIST security category 1, as per Equation˜1. Ternary-SDP and CROSS-SDP are parametrized according to [23] and [21].
TABLE II: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the VOLEitH framework. The proof system parameters are selected analogously to those in SDitH [3] to ensure comparable runtime performance of the VC scheme.
Security category $\mathcal{E}$-SDP Proof System Performance
[…]
NIST 1 $127$ 77 $127$ $76$ $16$ $256$ $32$ $101$ 33 fast $6432$
$11$ $2048$ $22$ $107$ $10$ short $4372$
33 22 $579$ $213$ $16$ $256$ $96$ $101$ 11 fast $3515$
$11$ $2048$ $81$ $107$ 88 short $2974$
NIST 3 $127$ 77 $187$ $111$ $24$ $256$ $48$ $153$ 33 fast 14 35914\,359
$16$ $4096$ $32$ $157$ 33 short $9361$
33 22 $839$ $309$ $24$ $256$ $144$ $153$ 11 fast $7816$
$16$ $4096$ $128$ $157$ 11 short $6463$
NIST 5 $127$ 77 $251$ $150$ $32$ $256$ $64$ $206$ 33 fast 25 57325\,573
2121 $4096$ $42$ $216$ 77 short 16 58916\,589
33 22 $1102$ $406$ $32$ $256$ $192$ $206$ 11 fast 13 85113\,851
2121 $4096$ $168$ $216$ 55 short 11 52111\,521
IV Concrete Performance Evaluation
Performance TCitH. From the provided modeling of $\mathcal{E}$-SDP, a digital signature is derived using the TCitH framework [8] and the Fiat-Shamir Transform, as summarized in Section˜II. The chosen parameters and the obtained signature sizes are provided in Table˜I. The parameters $\tau$, $N$, and $w$ of the proof system are selected following [6] to allow for comparable runtime performance of the VC scheme. To simplify implementation, the smallest extension degree $\mu=[\mathbb{K}:\mathbb{F}_{p}]$ satisfying the requirement $N\leq\left\lvert\mathbb{K}\right\rvert=\left\lvert\mathbb{F}_{p}\right\rvert^{\mu}$ is chosen. This determines the batching parameter $\eta$.
@@ IV Concrete Performance Evaluation
Performance VOLEitH. Table˜II provides an overview of the performances that can be achieved if TCitH is replaced by the VOLEitH framework. Parameters of the proof system are chosen analogously to SDitH [3]. For Ternary-SDP, only a minor improvement in signature size is achievable. A similar effect is observed in [6]: for OWFs with small degree and small witness size, the structurally simpler TCitH framework remains competitive. On the other hand, for CROSS-SDP, which has a degree of $d=z=7$, VOLEitH enables a reduction in signature size of about $1\text{\,}\mathrm{kB}$ for NIST category 1, and of up to $6\text{\,}\mathrm{kB}$ for NIST category 5.
Comparison with NIST submissions. A comparison of the achieved signature sizes with other ZK-based signatures submissions to NIST’s call for additional digital signature schemes [25] is provided in Table˜III. As can be seen, the TCitH and VOLEitH frameworks can be combined with a wide range of OWFs. The proposed modeling of $\mathcal{E}$-SDP is competitive compared to the state-of-the-art: for the Ternary-SDP, we obtain a signature size which is very close to those of the smallest proposals that utilize the TCitH or the VOLEitH framework with an alternative hardness assumption. The sizes obtained from CROSS-SDP are slightly larger than those of SDitH [3] and PERK [4], which also use modelings of degree greater than two, but smaller than $z=7$. The code-based proposals CROSS and LESS target other trade-offs: CROSS uses a simple CVE-like protocol [26, 12], and LESS achieves a small signature size by using multiple public keys (resulting in a public-key size that is multiple times larger than those of the other schemes shown) [11]. SQISign [15] is significantly more compact than the other schemes; this comes at the cost of a complex and slow construction.
TABLE III: Comparison with other ZK-based signatures. Shown sizes are for the short optimization tradeoff; runtime may vary.
Scheme Assumption Design $\left\lvert\sigma\right\rvert$ in $\mathrm{kB}$]
@@
SQISign Supersingular Isogenies sigma protocol $0.15$
Comparison with [27]. Recently [27] proposed a new MPCitH-based protocol for $\mathcal{E}$-SDP. Instead of relying on the TCitH and VOLEitH frameworks, their construction is built from the ground up, designing an MPC protocol that computes the syndrome corresponding to a shared restricted vector and turning it into a digital signature using the classical MPCitH framework. The approach presented in this paper offers greater flexibility in parameter selection and achieves smaller signatures, in particular through Ternary-SDP.
V Conclusion